# Long-Lived Photon Blockade with Weak Optical Nonlinearity

### v1

@@ Abstract
Introduction—Photon blockade is a quantum effect whereby one photon in a nonlinear resonator blocks the entry of other photons, giving rise to antibunched photon statistics [1, 2, 3, 4]. Conventionally, photon blockade requires the optical nonlinearity within a cavity to exceed its loss: i.e., $\alpha\gg\gamma$, where $\alpha$ characterizes the strength of the nonlinearity and $\gamma$ is the cavity decay rate. This can be interpreted in terms of a two-photon state being driven off-resonance by the nonlinearity, shifting its energy by more than the cavity linewidth [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Achieving this in the optical regime (including telecom wavelengths) is challenging, as material nonlinearities tend to be very weak relative to cavity linewidths. Systems featuring exceptionally strong optical nonlinearities are required, such as exciton-polaritons [24, 25], trions [26], Rydberg excitons [27], and Moire excitons [28].
The strong-nonlinearity condition may be bypassed via the phenomenon of unconventional photon blockade (UPB), which uses destructive interference between two or more nonlinear cavities (or cavity modes) to cancel the two-photon amplitudes [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50]. It is commonly believed that UPB requires the inter-cavity coupling rate $J$ to satisfy [29]
$J\gg\gamma.$ (1)
This seems to make intuitive sense: for interference to play a role, the photons must be able to hop many times before leaking away. In this regime, it then turns out that even a weak nonlinearity, $\alpha\ll\gamma$, can produce photon antibunching. However, even if the second-order photon correlation $g^{(2)}(\tau)$ vanishes at $\tau=0$, the suppression only holds over a time window of
@@
We assume a single-input, single-output setup where a cavity $d$ is driven, and we look at the photon populations in a “readout” cavity (which may or may not be $d$). The driving Hamiltonian is $\mathcal{H}_{d}=F_{d}a_{d}^{\dagger}+F^{*}_{d}a_{d}$, where $F_{d}$ is the complex driving amplitude. (The following derivation can also be straightforwardly generalized to the case of multiple coherently driven cavities.) The density matrix $\rho$ is then governed by the Lindblad master equation [52]
@@
We seek a solution that is stationary (i.e., co-rotating with the drive, $d/dt\rightarrow 0$). We truncate to superpositions of only 1- and 2-photon states, and derive the following expression for the equal-time second-order correlation function:
@@
where $|i\rangle=a_{i}^{\dagger}|\text{vac}\rangle$ denotes the state with one photon in cavity $i$ and $|i,j\rangle\equiv|i\rangle\otimes|j\rangle$. This generalizes a result reported by us for a specific lattice model [50]; the details of the derivation are given in the Supplemental Materials [53]. In Eq. (8), the expectation values are taken over the aforementioned stationary state, while in Eq. (9) $|\bar{\psi}^{(1)}\rangle$ and $|\bar{\psi}^{(2)}\rangle$ denote projections of the steady-state wavefunction to the 1- and 2-photon subspaces respectively.
Using Eq. (9), we can construct the following account of how photon antibunching arises. By fine-tuning the system parameters, it may be possible to satisfy $\langle i|\bar{\psi}^{(1)}\rangle=0$, which we call a 1-photon dark state, for some cavity $i$. In the absence of nonlinearity, such a condition also implies a 2-photon dark state $\langle i,j|\bar{\psi}^{(2)}\rangle=0$, since $|\bar{\psi}^{(2)}\rangle=|\bar{\psi}^{(1)}\rangle\otimes|\bar{\psi}^{(1)}\rangle/\sqrt{2}$. We then find that even though the numerator and denominator of (9) both vanish, $f_{ij}(z)$ does not vanish; in fact, $g_{ij}^{(2)}=1$ for all $z$. A weak nonlinearity, however, can induce a small misalignment between the 1-photon and 2-photon dark states. Given a 1-photon dark state at $z_{0}$, the 2-photon dark state condition is shifted to a nearby value of $z_{0}+\delta z$. At this point, the numerator of (9) vanishes while the denominator is nonzero. This is our desired photon antibunching point.
@@
UPB with two cavities—For a two-cavity system, with $J_{12}=J_{21}=J\in\mathbb{R}^{+}$, the coupling matrix has eigenvalues $\epsilon_{1,2}=\pm J$ and the eigenbasis $|\varphi_{1,2}\rangle=2^{-1/2}(|1\rangle\pm|2\rangle)$. Eqs. (10)–(12) predict antibunching in the driven cavity ($d=1$), $g_{11}^{(2)}(0)=0$, if and only if $z^{3}\approx-\alpha J^{2}/2$ [53]. The three roots correspond to three 2-photon dark states, and are depicted by blue circles in Fig. 1(a). They lie close to the 1-photon dark state at $z=0$, for which $\langle 1|\bar{\psi}^{(1)}\rangle\propto G_{11}=z(z^{2}-J^{2})^{-1}=0$, as indicated by the red cross in the figure. We focus on the root with positive cavity loss, for which
@@
The fast oscillation in $\cos(J\tau)$ yields a small time window of $\delta\tau\sim 1/J$, tied to the large coupling eigenvalues ±$\pm J$.
Photon antibunching in this model follows the mechanism laid out by Eqs. (8)–(12). The 2-photon dark state is perturbed away from the 1-photon dark state, and since the latter occurs at $z=0$, the solution we arrive at involves small $\gamma$ (relative to $J$). The lattice model studied in Ref. 50 also works similarly.
Figure 1: (a) Schematic of the zeros (circles) and pole (cross) of $f_{11}(z)$ in the complex $z$ plane, for a two-cavity model. The pole occurs at $z=0$, and UPB occurs at the zeros. (b) Schematic of a four-cavity setup that can exhibit long-lived photon blockade (LLPB). (c) Distribution of zeros (circles) and pole (cross) of $f_{22}(z)$ for the four-cavity model of (b). This illustration is not drawn to scale. (d) Equal-time two-photon correlation function $g^{(2)}(0)$ for the four-cavity system versus cavity detuning $\Delta$ and loss rate $\gamma$. Upper plot: analytical results obtained from Eqs. (20)–(22). Near the pole (central red spot), there are two zeros of $g^{(2)}(0)$ (blue spots) corresponding to LLPB. The model parameters are $k=16$, $J=0.1227$, $J^{\prime}=0.02454$, and $\alpha=0.001227$, chosen so that one of the LLPB points occurs at $\gamma=1$. Lower plot: WFMC simulation results, obtained by directly solving Lindblad master equation with the same parameters and $F_{d}=10^{-5}$.
Long-Lived Photon Blockade—The above scenario can be bypassed if we find a configuration of cavities with a 1-photon dark state at large $\gamma$. Specifically, we want the linear single-photon Green’s function to satisfy $G_{id}(z)=0$ for some $i,d$, with $\gamma=-2\mathrm{Im}(z)\gg\mathrm{max}\,|J_{ij}|$. This turns out to be impossible for $N=2$ and $N=3$ (though if we relax the assumption that the cavities are identical, we can find a zero for $N=3$ if one cavity is completely lossless); for details, see the Supplemental Materials [53]. For $N=4$, however, the condition can be satisfied.
Consider the configuration shown in Fig. 1(b), with four cavities arranged in a ring with couplings $J_{12},J_{23},J_{34},J_{41}\in\mathbb{R}^{+}$; the driven cavity is $d=1$, and photons are extracted from cavity $i=2$. A Dyson series expansion in the weak-coupling regime [54] indicates that $G_{12}(z)=G_{21}(z)$ can have a zero that meets the above conditions if $J_{12}\ll J_{23},J_{34},J_{41}$. Specifically, we consider
@@
LLPB is predicted to occur at $z\approx z_{0}\pm\delta z$, as shown schematically in Fig. 1(c).
In the upper plot of Fig. 1(d), we plot $g_{22}^{(2)}(0)$ versus the cavity parameters $\Delta$ and $\gamma$ using Eqs. (20)–(22). The parameters, stated in the figure caption, satisfy $\alpha\ll\gamma$ (weak nonlinearity) and $J^{\prime}/k\ll J^{\prime}\ll J\ll\gamma$ (weak coupling). For comparison, the lower plot of Fig. 1(d) shows the results obtained based on wavefunction Monte Carlo (WFMC) [55, 56, 57, 58] simulations, which directly solve the Lindblad master equation (4) using $F_{d}=10^{-5}$, with all other parameters kept the same. It is worth noting that the WFMC simulations give $g_{22}^{(2)}(0)\ll 1$, but not exactly zero, in the weak-drive regime. This value of $g_{22}^{(2)}(0)$ increases with $F_{d}$, alongside an increase in the mean photon occupation number in the signal cavity [53].
@@
Conclusion—We have re-analyzed the origins of unconventional photon blockade (UPB), in which photon antibunching is achieved by quantum interference in weakly nonlinear coupled cavities [29]. A major problem with UPB has hitherto been the rapid oscillation in $g^{(2)}(\tau)$ due to strong inter-cavity coupling [37, 39]. We show that this can be overcome by designing a coupled-cavity system so that the linear single-photon Green’s function exhibits a zero (a 1-photon dark state) at a large cavity loss rate. This leads to a “long-lived photon blockade” (LLPB) scenario whereby $g^{(2)}(\tau)\propto\tau^{4}$ and the time window exceeds that of conventional photon blockade under strong nonlinearity. It is remarkable that shortening the cavity lifetime can effectively lengthen the antibunching time window, which can be regarded as a fresh example of non-Hermiticity (in the form of dissipation) producing unexpected physical behaviors [59, 60]. Our proposal does not need one-way inter-cavity couplings or other exotic requirements [51, 39], and should be highly compatible with quantum optics experiments.

### v2

@@ Abstract
Introduction—Photon blockade is a quantum effect whereby one photon in a nonlinear resonator blocks the entry of other photons, giving rise to antibunched photon statistics [1, 2, 3, 4]. Conventionally, photon blockade requires the optical nonlinearity within a cavity to exceed its loss: i.e., $\alpha\gg\gamma$, where $\alpha$ characterizes the strength of the nonlinearity and $\gamma$ is the cavity decay rate. This can be interpreted in terms of a two-photon state being driven off-resonance by the nonlinearity, shifting its energy by more than the cavity linewidth [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Achieving this in the optical regime (including telecom wavelengths) is challenging, as material nonlinearities tend to be very weak relative to cavity linewidths. Systems featuring exceptionally strong optical nonlinearities are required, such as exciton-polaritons [25, 26], trions [27], Rydberg excitons [28], and Moire excitons [29].
The strong-nonlinearity condition may be bypassed via the unconventional photon blockade (UPB) phenomenon, which uses destructive interference between two or more nonlinear cavities (or cavity modes) to cancel the two-photon amplitudes [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. It is commonly believed that UPB requires the inter-cavity coupling rate $J$ to satisfy
$J\gg\gamma,$ (1)
which seems to make intuitive sense: for interference to play a role, the photons must be able to hop many times before leaking away [30]. In this regime, it then turns out that even a weak nonlinearity, $\alpha\ll\gamma$, can produce photon antibunching. However, even if the second-order photon correlation $g^{(2)}(\tau)$ vanishes at $\tau=0$, the suppression only holds over a time window of
@@
We assume a single-input, single-output setup where a cavity $d$ is driven, and we look at the photon populations in a “readout” cavity (which may or may not be $d$). The driving Hamiltonian is $\mathcal{H}_{d}=F_{d}a_{d}^{\dagger}+F^{*}_{d}a_{d}$, where $F_{d}$ is the complex driving amplitude. (The following derivation can also be straightforwardly generalized to the case of multiple coherently driven cavities.) The density matrix $\rho$ is governed by the Lindblad master equation [54]
@@
We seek a solution that is stationary (i.e., co-rotating with the drive, $d/dt\rightarrow 0$). Truncating to superpositions of only 1- and 2-photon states yields the following equal-time second-order correlation function:
@@
where $|i\rangle=a_{i}^{\dagger}|\text{vac}\rangle$ denotes the state with one photon in cavity $i$ and $|i,j\rangle\equiv|i\rangle\otimes|j\rangle$. The full derivation, which generalizes a previous result based on a specific lattice model [52], is given in the Supplemental Materials [55]. In Eq. (8), the expectation values are taken over the stationary state, while in Eq. (9) $|\bar{\psi}^{(1)}\rangle$ and $|\bar{\psi}^{(2)}\rangle$ denote projections of the steady-state wavefunction to the 1- and 2-photon subspaces respectively.
Using Eq. (9), we construct the following account of how photon antibunching arises. By fine-tuning the system parameters, it may be possible to satisfy $\langle i|\bar{\psi}^{(1)}\rangle=0$, which we call a 1-photon dark state, for some cavity $i$. In the absence of nonlinearity, such a condition also implies a 2-photon dark state $\langle i,j|\bar{\psi}^{(2)}\rangle=0$, since $|\bar{\psi}^{(2)}\rangle=|\bar{\psi}^{(1)}\rangle\otimes|\bar{\psi}^{(1)}\rangle/\sqrt{2}$. We find that even though the numerator and denominator of (9) both vanish, $f_{ij}(z)$ does not vanish; in fact, $g_{ij}^{(2)}=1$ for all $z$. A weak nonlinearity, however, can induce a small misalignment between the 1-photon and 2-photon dark states. Given a 1-photon dark state at $z_{0}$, the 2-photon dark state condition is shifted to a nearby value of $z_{0}+\delta z$. At this point, the numerator of (9) vanishes while the denominator is nonzero. This is our desired photon antibunching point.
Figure 1: (a) Schematic of zeros (circles) and pole (cross) of $f_{11}(z)$ in the complex $z$ plane, for a two-cavity model. UPB occurs at the zeros. (b) Schematic of a four-cavity system exhibiting long-lived photon blockade (LLPB). (c) Schematic of a photonic structure that can implement the model in (b), using two coupled microring resonators with clockwise and counterclockwise modes of each ring. (d) Distribution of zeros (circles) and pole (cross) of $f_{22}(z)$ for the four-cavity model of (b). This illustration is not drawn to scale. (e) Equal-time two-photon correlation function $g^{(2)}(0)$ for the four-cavity system versus cavity detuning $\Delta$ and loss rate $\gamma$. Upper plot: analytical results obtained from Eqs. (20)–(22). Near the pole (central red spot), there are two zeros of $g^{(2)}(0)$ (blue spots) corresponding to LLPB. The model parameters are $k=16$, $J=0.1227$, $J^{\prime}=0.02454$, and $\alpha=0.001227$, chosen so that one of the LLPB points occurs at $\gamma=1$. Lower plot: WFMC simulation results, obtained by directly solving Lindblad master equation with the same parameters and $F_{d}=10^{-5}$.
@@
UPB with two cavities—For a two-cavity system, with $J_{12}=J_{21}=J\in\mathbb{R}^{+}$, the coupling matrix has eigenvalues $\epsilon_{1,2}=\pm J$ and the eigenbasis $|\varphi_{1,2}\rangle=2^{-1/2}(|1\rangle\pm|2\rangle)$. Eqs. (10)–(12) predict antibunching in the driven cavity, $g_{11}^{(2)}(0)=0$, if and only if $z^{3}\approx-\alpha J^{2}/2$ [55]. The three roots, corresponding to three 2-photon dark states, are depicted by blue circles in Fig. 1(a). They lie close to the 1-photon dark state at $z=0$, for which $\langle 1|\bar{\psi}^{(1)}\rangle\propto G_{11}=z(z^{2}-J^{2})^{-1}=0$, marked by a red cross in the figure. We focus on the root with positive cavity loss, for which
@@
The fast oscillation in $\cos(J\tau)$ yields a small time window of $\delta\tau\sim 1/J$, tied to the large coupling eigenvalues ±$\pm J$. As laid out by Eqs. (8)–(12), photon antibunching occurs when the 2-photon dark state is perturbed away from the 1-photon dark state; since the latter occurs at $z=0$, the solution involves small $\gamma$ (relative to $J$). The lattice model studied in Ref. 52 also works similarly.
Long-Lived Photon Blockade—We now present a system with a 1-photon dark state at large $\gamma$. We want the linear single-photon Green’s function to satisfy $G_{id}(z)=0$ for some $i,d$, with $\gamma=-2\mathrm{Im}(z)\gg\mathrm{max}\,|J_{ij}|$, but this turns out to be impossible for $N=2$ and $N=3$ (if we relax the assumption that the cavities are identical, there can be a zero for $N=3$ if one cavity is completely lossless); for details, see the Supplemental Materials [55]. For $N=4$, however, the condition can be satisfied with the setup in Fig. 1(b), with four cavities arranged in a ring with couplings $J_{12},J_{23},J_{34},J_{41}\in\mathbb{R}^{+}$; the driven cavity is $d=1$, and photons are extracted from cavity $i=2$. This configuration can be implemented using photonic structures like the one shown in Fig. 1(c), with two ring resonators each hosting a pair of weakly-coupled clockwise and counterclockwise modes [56, 57, 58, 59]. In the Supplemental Materials, we present further details about this exemplary coupled-ring setup, including full-wave simulations showing how it can achieve the desired model parameters [55].
A Dyson series expansion in the weak-coupling regime [60] indicates that $G_{12}(z)=G_{21}(z)$ can have a zero that meets the above conditions if $J_{12}\ll J_{23},J_{34},J_{41}$. Specifically, we consider
@@
Therefore, LLPB occurs at $z\approx z_{0}+\delta z$, as shown schematically in Fig. 1(d).
In the upper panel of Fig. 1(e), we plot $g_{22}^{(2)}(0)$ versus the cavity parameters $\Delta$ and $\gamma$ using Eqs. (20)–(22). The parameters, stated in the figure caption, satisfy $\alpha\ll\gamma$ (weak nonlinearity) and $J^{\prime}/k\ll J^{\prime}\ll J\ll\gamma$ (weak coupling). The lower panel of Fig. 1(e) shows results from wavefunction Monte Carlo (WFMC) [61, 62, 63, 64] simulations that directly solve the Lindblad master equation (4), using $F_{d}=10^{-5}$ with all other parameters kept the same. Note that the WFMC simulations give $g_{22}^{(2)}(0)\ll 1$, but not exactly zero, in the weak-drive regime; the value of $g_{22}^{(2)}(0)$ increases with $F_{d}$, alongside an increase in the mean photon occupation number in the signal cavity [55].
@@
Conclusion—We have re-analyzed the origins of unconventional photon blockade (UPB), in which photon antibunching is achieved by quantum interference in weakly nonlinear coupled cavities [30]. A major problem with UPB has hitherto been the rapid oscillation in $g^{(2)}(\tau)$ due to strong inter-cavity coupling [38, 41]. We show that this can be overcome by designing a coupled-cavity system so that the linear single-photon Green’s function exhibits a zero (a 1-photon dark state) at a large cavity loss rate. This leads to a “long-lived photon blockade” (LLPB) scenario whereby $g^{(2)}(\tau)\propto\tau^{4}$ and the time window exceeds that of conventional photon blockade under strong nonlinearity. It is remarkable that shortening the cavity lifetime can effectively lengthen the antibunching time window, which can be regarded as a fresh example of non-Hermiticity (in the form of dissipation) producing unexpected physical behaviors [65, 66]. Our proposal does not need one-way inter-cavity couplings or other exotic requirements [53, 41], and should be compatible with quantum optics experiments using setups like the one shown in Fig. 1(c) [55].
@@
S6 Photonics Realization
In this section, we explore how the quantum optical model could be implemented in practice. To achieve the four coupled optical modes depicted in Fig. 1(b) of the main text, we can employ either four coupled microring resonators, as shown in Fig. S5(a), or two coupled microring resonators with clockwise (CW) and counter-clockwise (CCW) modes, as shown in Fig. S5(b) and Fig. 1(c) of the main text. We will focus on the latter case.
Figure S5: Photonic implementations of the model. (a) Configuration using four coupled microring resonators. (b) Configuration using two coupled microring resonators where the clockwise (CW) and counter-clockwise (CCW) modes within each ring are coupled. (c) Inter-ring coupling strength $J$ as a function of the gap size $d_{gap}$ between the two rings, calculated using Finite Element Method (FEM) simulations.
In the two-ring setup, the CW and CCW modes have the same natural frequencies. They are coupled to each other, with coupling rates $J^{\prime}$ and $J^{\prime\prime}$, which can be achieved by deliberately-introduced subwavelength defects or surface roughness, as shown in previous studies [56, 59, 58, 57]. Assuming the two rings have the same shape and size, the Langevin equation for the system is given by
$i\frac{d}{dt}\begin{pmatrix}a_{1}\\ a_{2}\\ a_{3}\\ a_{4}\end{pmatrix}=\begin{pmatrix}\Delta-\frac{i}{2}(\gamma+\gamma_{in}+\gamma_{out})&J^{\prime\prime}&0&J\\ J^{\prime\prime}&\Delta-\frac{i}{2}(\gamma+\gamma_{in}+\gamma_{out})&J&0\\ 0&J&\Delta-\frac{i}{2}\gamma&J^{\prime}\\ J&0&J^{\prime}&\Delta-\frac{i}{2}\gamma\end{pmatrix}\begin{pmatrix}a_{1}\\ a_{2}\\ a_{3}\\ a_{4}\end{pmatrix}+\begin{pmatrix}i\sqrt{\gamma_{\textrm{in}}}\alpha_{\textrm{in}}\\ 0\\ 0\\ 0\end{pmatrix},$ (S59)
where $\alpha_{\textrm{in}}$ is the input field amplitude, $\gamma$ is the intrinsic decay rate of the rings, $\gamma_{\textrm{in}}$ is the input coupling rate, and $\gamma_{\textrm{out}}$ is the output coupling rate. The output field is given by $\alpha_{\textrm{out}}=\sqrt{\gamma_{\textrm{out}}}a_{2}$. The parameters $J$, $J^{\prime}$, and $J^{\prime\prime}$ represent the coupling strength between the two rings, the coupling strength between the CW and CCW modes within the second ring, and the coupling strength between the CW and CCW modes within the first ring, respectively. The detuning $\Delta=\omega_{c}-\omega_{L}$ is the difference between the cavity resonance frequency $\omega_{c}$ and the laser frequency $\omega_{L}$. (We set $\hbar=1$.)
The Hamiltonian terms describing the self-Kerr and cross-Kerr nonlinearities (between CW and CCW modes) are
$\mathcal{H}_{nl}=\sum_{i}\alpha a_{i}^{\dagger}a_{i}^{\dagger}a_{i}a_{i}+2\alpha a_{1}^{\dagger}a_{1}a_{2}^{\dagger}a_{2}+2\alpha a_{3}^{\dagger}a_{3}a_{4}^{\dagger}a_{4},$ (S60)
where $\alpha$ is the Kerr coefficient of each ring resonator.
We perform Finite Element Method (FEM) simulations of this setup, using exemplary parameters drawn from the experimental literature. We assume each ring is composed of silicon carbide (SiC), with radius $R=3\,\mu$m, waveguide thickness $t=0.35$ $\mu$m, and waveguide width $w=0.8\,\mu$m, operating at a wavelength near $\lambda=1550$ nm. The refractive indices are $n_{\textrm{SiC}}=2.45$ for the waveguides and $n(\text{SiO}_{2})=1.44$ for the cladding. The nonlinear refractive index for SiC is $n_{2}=4.8\times 10^{-6}$ $\mu$m2/W. The nonlinearity parameter $\alpha$ is estimated by [11, 38]
$\alpha=\frac{c\hbar^{2}\omega_{c}^{2}}{n_{\textrm{SiC}}^{2}V}n_{2},$ (S61)
where $V$ is the mode volume, approximated here by the volume of the microring resonator $V=2\pi Rwt$. This yields $\alpha\approx 4.7\times 10^{-6}$ $\mu$eV.
Figure S5(c) shows the coupling strength $J$ as a function of the gap size $d_{\textrm{gap}}$ between the two rings, derived from FEM simulations. In these simulations, the eigenenergy splitting between the symmetric and antisymmetric supermodes of the coupled system corresponds to $2J$, which is the eigenvalue difference of the matrix
$\begin{pmatrix}\Delta&J\\ J&\Delta\end{pmatrix}.$
The FEM simulation identifies an eigenmode at $\lambda=1553$ nm with a simulated intrinsic quality factor $Q_{\textrm{int}}\approx 1.5\times 10^{6}$. Accounting for other potential loss sources, we estimate the total intrinsic loss rate of the cavity as $\gamma=5$ $\mu$eV, corresponding to an unloaded quality factor $Q_{\textrm{cav}}=\hbar\omega_{c}/\gamma\approx 1.6\times 10^{5}$. The input and output coupling rates are chosen to be $\gamma_{\textrm{in}}=\gamma_{\textrm{out}}=2.5$ $\mu$eV. The couplings strengths $J^{\prime}$ and $J^{\prime\prime}$, can be controlled either by adjusting the gap size $d_{\textrm{gap}}$ in the four-cavity configuration [Fig. S5(a)] or by introducing a subwavelegnth defect or slightly increase surface roughness [56, 58, 57, 59], that weakly couple the CW and CCW modes in each ring for the two-cavity setup (Fig. S5(b)). To achieve LLPB, we require $J^{\prime\prime}\ll J^{\prime}$, but their actual values do not need to be precisely controlled; thus, we take $J^{\prime}=0.2$ $\mu$eV and $J^{\prime\prime}=0.02$ $\mu$eV.
Since experimental parameters may deviate from these nominal values, the detuning $\Delta=\hbar(\omega_{c}-\omega_{L})$ and the inter-ring coupling $J$ are treated as tunable parameters. The detuning $\Delta$ can be adjusted by tuning the laser frequency $\omega_{L}$. Fine-tuning the coupling $J$ can be achieved by first selecting a gap size $d_{\textrm{gap}}$ corresponding to the desired approximate value of $J$ (from Fig. S5(c)) and then continuously adjusting it. This fine-tuning can be accomplished by illuminating the gap region with a control laser source. Modulating the intensity of this control source allows for precise regional adjustments of the refractive index. Figure S6(b) illustrates the dependence of $J$ on the regional refractive index change $\Delta n$ within a spot of radius $r=0.5$ $\mu$m located in the gap (schematically shown in Fig. S6(a)), assuming an initial gap size of $d_{\textrm{gap}}=1.74$ $\mu$m.
Figure S6(c) displays a colormap of the zero-delay second-order correlation function $g^{(2)}(0)$ as a function of the detuning $\Delta$ and coupling strength $J$. The black contour line indicates the boundary of the antibunching region, defined by $g^{(2)}(0)<0.5$. The inset shows the time-delayed second-order correlation function $g^{(2)}(\tau)$ for the output field, calculated at the parameter values marked by the purple cross in the main panel. It shows a over $1$ ns time window for $g^{(2)}(\tau)<0.5$.
Figure S6: Fine-tuning and simulation results. (a) Schematic illustrating the $r=0.5$ $\mu$m spot within the inter-ring gap used for refractive index-based fine-tuning of the coupling $J$. (b) Dependence of the coupling strength $J$ on the regional refractive index change $\Delta n$ applied to the spot shown in (a), assuming a fixed gap size $d_{\textrm{gap}}=1.74$ $\mu$m. (c) Colormap of the zero-delay second-order correlation function $g^{(2)}(0)$ for the output field as a function of detuning $\Delta$ and coupling strength $J$. The black contour line ($g^{(2)}(0)=0.5$) delineates the antibunching region. The inset displays the time-delayed second-order correlation $g^{(2)}(\tau)$ calculated at the parameter point indicated by the purple cross.